OpenStudy (anonymous):

HELP! The force of attraction between two magnets F Newtons is inversely proportional to the square of the distance, d between the magnets. When the magnets are certain distance, the force is 10 Newtons. What is the force when this distance is doubled? -Please show your working as well.

OpenStudy (anonymous):

\[F=k/d ^{2}\]

OpenStudy (anonymous):

F=10 N, when distance is doubled, the distance squared is multiplied by 4 so the force would be 1/4. So the answer is F=10/4 Newtons or 2.5 Newtons

OpenStudy (anonymous):

Think of it this way:\[\frac{1}{d^2}\longrightarrow \frac{1}{(2d)^2}=\frac{1}{4}\left(\frac{1}{d^2}\right)\], which is clearly one quarter of the original.

OpenStudy (anonymous):

So I multiply the force 10N with 1/4 as well?

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

Okay, thank you so much! :)

OpenStudy (anonymous):

nice dude yakeyglee